OpenStudy (anonymous):

If $240 is invested at an interest rate of 9% per year and is compounded monthly, how much will the investment be worth in 14 years?

OpenStudy (anonymous):

use a calculator

OpenStudy (anonymous):

thee formula is A=Pe^rt

OpenStudy (anonymous):

P is the initial investment e is a button on the calculator r is the interest rate and t is the time

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

Okay I think I got now thank you so much!

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

says compounded monthy, not continuously

OpenStudy (anonymous):

yeah but its the same equation

OpenStudy (anonymous):

\[240(1+\frac{.09}{12})^{12\times 14}\] is the formula you need

OpenStudy (anonymous):

oh you are right im sorry

OpenStudy (anonymous):

no, monthy is monthly, continuously is continuously they are close, but once a month is not the same as continuous

OpenStudy (phi):

Did you get the answer?

OpenStudy (phi):

You can type this into the google search window 240*(1+0.09/12)^(12*14)= to get the number. Or use a calculator.